A father is 4 time as old as his son if their total age is 60 years, how old is each?
Answer by addingup(3677)

F = 4S father is 4 times son
F+S = 60 substitute for F:
4S+S = 60
5S = 60
S = 12 The son is 12. And the father:
12*4 = 48 years old is the father
:
J